OpenStudy (anonymous):

Two cards are drawn from a standard deck of cards. If they are drawn with replacement, what is the probability that both cards are 7s?

OpenStudy (aveline):

How many cards are in a deck? How many 7's are there in a deck?

OpenStudy (anonymous):

\[\left(\frac{4}{52}\right)^2=\frac{1}{169}=0.005917 \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome.